# Dynamics of Meta-learning Representation in the Teacher-student Scenario

### v1

@@ III.5 Disorder Averages
where the detailed expressions can be found in Ref. [40]. Employing the above results, we eventually conclude that the equation of motion of the averaged order parameters $Q_{kl}$ is
Finally, the meta-generalization error can be computed as
Eqs. (38), (40) and (41) are the major theoretical results of this work. Solving the ordinary differential equations Eqs. (38) and (40) with initial conditions for $\{R_{kn},Q_{kl}\}$ at $\alpha=0$, we obtain a macroscopic description of the learning dynamics through the order parameters.
IV Results
IV.1 Comparison to Simulation
Before analyzing various learning scenarios, we compare the theory of meta-learning dynamics developed in Sec. III to numerical experiments, as a sanity check of the theoretical derivation. For this purpose, we adopt the same task and data generation processes as in Sec. III.1 to conduct numerical simulations. The representation mapping of the meta-teacher network is generated as $\boldsymbol{B}_{n}\sim\mathcal{N}(0,\frac{1}{N}I_{N})$. At time $\alpha=0$, the representation mapping of the meta-learner is also initialized as $\boldsymbol{J}_{k}\sim\mathcal{N}(0,\frac{1}{N}I_{N})$. Note that $\boldsymbol{J}_{k}(\alpha=0)$ and $\boldsymbol{B}_{n}$ are statistically independent. The meta-learner is then trained using the FO-ANIL algorithm on the tasks generated by the meta-teacher. In this way, the quantities $R_{kn}$ and $Q_{kl}$ at the initial time corresponding to the simulation setup can be computed; they will be set as the initial conditions of the ordinary differential equations Eqs. (38) and (40), which remain to be solved to obtain the theoretical predictions.
The other system parameters are set as $N=1000,K=M=3,P=V=100,\eta_{J}=6,\eta_{w}=4$. The results for numerical simulations and the corresponding theoretical predictions are shown in Fig. 2. It can be observed that the theoretical curves of $\epsilon_{g}^{\text{meta}}$ and $\{R_{kn}\}$ closely matches the experimental data, which suggests that the theoretical model effectively captures the underlying phenomena observed in the experiments. The meta-generalization error $\epsilon_{g}^{\text{meta}}$ decrease as the training process evolves, suggesting that the meta-learner effectively learns the knowledge of the meta-teacher. On the other hand, the deviation of theory and simulation for $\{Q_{kl}\}$ is a bit more prominent for the chosen set of system parameters. In Appendix B, we provide an argument and further numerical results to support that the order parameters $\{Q_{kl}\}$ are subject to larger statistical error when $V$ is finite and $\eta_{J}$ and $\eta_{w}$ are large. When $P,V$ are sufficiently large, we expect that the theory provides an accurate description of the dynamics of both sets of order parameters $\{R_{kn}\}$ and $\{Q_{kl}\}$.
Figure 2: Comparison between theoretical predictions (marked by “TH”) and simulated experiments (marked by “Exp”) of the online meta-learning dynamics under consideration. The simulation set-up is outlined in Sec. IV.1, with system parameters $N=1000,K=M=3,P=V=100,\eta_{J}=6,\eta_{w}=4$. The simulations are averaged over $10$ independent runs, each with different realizations of training tasks and datasets, while maintaining the same initial conditions. The error bars represent one standard deviation across these $10$ trials. (a) The meta-generalization error $\epsilon_{g}^{\text{meta}}$ as a function of the normalized task index $\alpha=t/N$. Since directly evaluating $\epsilon_{g}^{\text{meta}}$ by testing on a large dataset is time-consuming, we instead prepare a small test dataset at each time step and compute a moving average of $\epsilon_{g}^{\text{meta}}$ over time to smooth out fluctuations for each trial, with a sliding window $\Delta\alpha=0.05$. Panels (b) and (c) depict the dynamic behavior of $Q$, where solid lines indicate experimental observations and dashed lines show theoretical results. Panels (d), (e), and (f) illustrate the dynamical evolution of $R$, with solid lines representing experimental data and dashed lines corresponding to theoretical predictions.
IV.2 Meta-Representation Learning
A sufficient condition for successful meta-learning is that the meta-learner’s representation mapping $\{\boldsymbol{J}_{k}\}$ aligns with the meta-teacher’s representation mapping $\{\boldsymbol{B}_{n}\}$, up to the permutation and reflection symmetries of the hidden units. [^2] By reflection symmetry, we mean that the solution of $\boldsymbol{J}_{k}$ is allow to aligned with the opposite direction of $\boldsymbol{B}_{n}$ for some hidden teacher neuron $n$ (suppose that $\boldsymbol{J}_{k}=-\boldsymbol{B}_{n}$); this is because the chosen activation function $g(x)=\operatorname{erf}(x/\sqrt{2})$ is an odd function, so that setting $w^{t}_{k}=-u^{t}_{n}$ enables the $t$-th learner to learn the output component contributed by the $n$-th hidden unit of the $t$-th teacher. This can be measured by monitoring the overlaps $\{R_{kn}:=\boldsymbol{J}_{k}^{\top}\cdot\boldsymbol{B}_{n}\}$ between the meta-learners’ and meta-teachers’ hidden units, and their cosine similarity defined by
Suppose that a successful alignment occurs with $R_{kn}\propto\delta_{k,n}$ (i.e., $\boldsymbol{J}_{k}\propto\boldsymbol{B}_{n}$); then, based on Eqs. (31) and (33), the $k$-th hidden-to-output weight of the $t$-th learner will, on average, be determined by the $n$-th hidden-to-output weight of the $t$-th teacher (i.e., $\langle w^{t}_{k}\rangle\propto u^{t}_{n}$). If the learning rate $\eta_{w}$ is properly set and the order parameters $\{Q_{kk}\}$ are accurately learned, the $t$-th learner could potentially reproduce the exact output of the $t$-th teacher. Consequently, we will focus on the similarity between the representation mappings $\mathbf{J}$ and $\mathbf{B}$ and examine their relationship with meta-generalization performance.
Inspired by [40], we set the system parameters as $K=M=3,\eta_{J}=3,T_{mn}=m\delta_{m,n}$, and consider the initial conditions $Q_{kl}=\frac{1}{2}\delta_{kl},R_{kn}=10^{-12}$. The results for two different choices of $\eta_{w}$ are shown in Fig. 3. From Fig. 3(a), we observe that $\epsilon_{g}^{\text{meta}}$ decreases with increasing $\alpha$ for both cases, indicating the improvement of the meta-generalization ability of the meta-learner.
For $\eta_{w}=3$, $\epsilon_{g}^{\text{meta}}$ experiences an initial drop and remains on a plateau for the duration of the analysis, up to $\alpha=500$. The initial drop occurs at around $\alpha=180$, at which time $R_{13}$, $R_{23}$, and $R_{33}$ (or $\rho_{13}$, $\rho_{23}$, and $\rho_{33}$) starts to grow from zero to a similar value as observed in Fig. 3(b) and (c). This implies that all the meta-learner’s hidden units align with only one hidden unit ($n=3$) of the meta-teacher network (as illustrated in Fig. 3(d)), which is a suboptimal solution.
On the other hand, for $\eta_{w}=9$, $\epsilon_{g}^{\text{meta}}$ undergoes three distinct drops, ultimately decreasing to a value close to zero. Analyzing the behaviors of $R_{kn}$ and Fig. 3(e), we observe that $R_{13}$, $R_{32}$, and $R_{21}$ progressively develop and eventually converge to distinct non-zero values, while the remaining order parameters $\{R_{kl}\}$ approach zero. A similar pattern is seen for the cosine similarities $\{\rho_{kl}\}$ shown in Fig. 3(f). At around $\alpha=40$, the meta-learner gets into the symmetric suboptimal state where all its hidden units learn the single hidden unit ($n=3$) of the meta-teacher, indicated by the growth of $R_{13},R_{23}$ and $R_{33}$. But eventually, the meta-learner is able to escape this symmetric state by further differentiating its hidden units. In the final stage, the hidden units in the meta-learner specialized to different hidden units in the meta-teacher network, i.e., $\boldsymbol{J}_{1}\propto\boldsymbol{B}_{3},\boldsymbol{J}_{2}\propto-\boldsymbol{B}_{1},\boldsymbol{J}_{3}\propto-\boldsymbol{B}_{2}$ as illustrated in Fig. 3(g). Up to the permutation and reflection symmetries, the meta-learner’s representation mapping perfectly align with that of the meta-teacher. Consequently, the meta-generalization error can ultimately be reduced to a value close to zero.
Figure 3: A case study of meta-representation learning. The system parameters are $K=M=3,\eta_{J}=3,T_{mn}=m\delta_{m,n}$, and the initial conditions are $Q_{kl}=\frac{1}{2}\delta_{k,l},R_{kn}=10^{-12}$. (a) $\epsilon_{g}^{\text{meta}}$ v.s. $\alpha$. (b) $R_{kn}$ v.s. $\alpha$ for $\eta_{w}=3$. (c) $\rho_{kn}$ v.s. $\alpha$ for $\eta_{w}=3$. (d) Pictorial illustration of the outcome of the meta-representation learning for $\eta_{w}=3$. (e) $R_{kn}$ v.s. $\alpha$ for $\eta_{w}=9$. (f) $\rho_{kn}$ v.s. $\alpha$ for $\eta_{w}=9$. (g) Pictorial illustration of the outcome of the meta-representation learning for $\eta_{w}=9$.
IV.3 The Role of Learning Rates
As seen in Sec. IV.2, it is crucial to determine the appropriate learning rate $\eta_{w}$ for the effectiveness of meta-learning. We expect that the choice of $\eta_{J}$ also matters. Here, we study the effects of learning rates in greater detail within our theoretical framework.
We set the same system parameters and initial conditions as those in the case study in Sec. IV.2, except that the learning rates $\eta_{J}$ and $\eta_{w}$ can vary. The meta-learning dynamics is evolved up to time $\alpha=500$, at which the final value of $\epsilon_{g}^{\text{meta}}$ is extracted for each choice of learning rates. The results are shown in Fig. 4. It can be observed from Fig. 4(a) that for a particular learning rate $\eta_{J}$, the final meta-generalization error first decreases rapidly and then increases gently when $\eta_{w}$ increases. There is an optimal value of $\eta_{w}$, but increasing it to a relatively large value does not deteriorate the performance a lot. In Fig. 4(b), we observe that for a particular learning rate $\eta_{w}$, the final meta-generalization error decreases and saturates when $\eta_{J}$ increases. The case with $\eta_{w}=6$ yields a better generalization performance than the case with $\eta_{w}=3$ for various choices of $\eta_{J}$.
Figure 4: Effect of learning rates for meta-generalization. The system parameters and initial conditions are the same as those in Sec. IV.2. (a) $\epsilon_{g}^{\text{meta}}$ v.s. $\eta_{w}$ when fixing $\eta_{J}$. (b) $\epsilon_{g}^{\text{meta}}$ v.s. $\eta_{J}$ when fixing $\eta_{w}$.
IV.4 Overparameterization
As $K$ increases, both the minimal values of $\eta_{w}$ and $\eta_{J}$ required for the meta-learner to achieve $\epsilon_{g}^{\text{meta}}=0.01$ before $\alpha_{\text{final}}=450$ also increase. This can be understood by referring to Eqs. 28, 29 and 31, which show that both $\eta_{w}$ and $\eta_{J}$ are scaled by the factor $\frac{1}{\sqrt{K}}$. Therefore, larger values of $\eta_{w}$ and $\eta_{J}$ are necessary to maintain a comparable pace of the dynamics when $K$ increases. More interestingly, as $K$ increases, the non-yellow region expands, indicating that in more overparameterized models, less stringent choices of the learning rates (particularly $\eta_{J}$) can be used to achieve a strong level of meta-generalization for a given number of tasks. This echos with previous findings that overparameterization facilitates optimization.
Figure 6: Meta-learning dynamics of an overparameterized model with $M=3,K=6$. The other system parameters and initial conditions are $\eta_{J}=3,\eta_{w}=9,T_{mn}=m\delta_{m,n},Q_{kl}=\frac{1}{2}\delta_{k,l},R_{kn}=10^{-12}$. (a) $\epsilon_{g}^{\text{meta}}$ v.s. $\alpha$. (b) $\rho_{kn}$ v.s. $\alpha$. It can be observed that $|\rho_{21}|,|\rho_{61}|$ converge to one, indicating that the 2nd and 6th hidden units in the meta-learner have specialized to the 1st hidden unit in the meta-teacher network. Similarly, $|\rho_{32}|,|\rho_{52}|$ converge to one, which suggests that the 3rd and 5th hidden units in the meta-learner have specialized to the 2nd hidden unit in the meta-teacher. Lastly, $|\rho_{13}|,|\rho_{43}|$ converge to one, which indicates that the 1st and 4th hidden units in the meta-learner have specialized to the 3rd hidden unit in the meta-teacher.
IV.5 Variability of Teacher Representation Mapping
In previous sections, we have primarily focused on the scenario that different task-specific teachers share the same representation mapping $\mathbf{B}$, derived from the meta-teacher network. Here, we relax this assumption by allowing task-specific teachers to have different representation mappings, while maintaining a certain level of similarity. To model this scenario, we define the input-to-hidden weights of the $t$-th teacher network as
where $\mathbf{B}$ is the common input-to-hidden weights of the meta-teacher shared across all tasks, while $\Delta\mathbf{B}^{t}$ is specific to the $t$-th task. The parameter $\gamma\in[0,1]$ controls the variability of the teacher weights across tasks. We assume that $\Delta\mathbf{B}^{t}$ is drawn from a certain distribution $p(\Delta\mathbf{B})$. When $\gamma$ is close to one, the contribution of the component $\sqrt{1-\gamma^{2}}\Delta\mathbf{B}^{t}$ is small, which allows us to perform a series expansion of the activation $g(\boldsymbol{B}_{m}^{t\top}\cdot\boldsymbol{\xi})=g(\gamma\boldsymbol{B}_{m}^{\top}\cdot\boldsymbol{\xi}+\sqrt{1-\gamma^{2}}\Delta\boldsymbol{B}_{m}^{t\top}\cdot\boldsymbol{\xi})$ around the value $\gamma\boldsymbol{B}_{m}^{\top}\cdot\boldsymbol{\xi}$. In this manner, we can derive the dynamical equations for the order parameters using the large $\gamma$ approximation. Further details of the derivation are given in Appendix. C.
Fig. 7(a) illustrates the evolution of $\epsilon_{g}^{\text{meta}}$ during the meta-training process for various values of $\gamma$. It demonstrates that even when $\gamma<1$, indicating variability in the teacher’s representation mapping, the FO-ANIL algorithm is still capable of effectively reducing the meta-generalization error, though not as optimally as in the case with $\gamma=1$. Fig. 7(b) depicts the effect of $\eta_{w}$ on $\epsilon_{g}^{\text{meta}}$ after a certain time window, which reassures the efficacy of the meta-learning algorithm for different choices of the learning rates. Similar to the findings in Sec. IV.3, there is an optimal value of $\eta_{w}$ for each case, but increasing it to a relatively large value does not significantly impair performance.
V Conclusion
For a fixed number of tasks (i.e., within a specific time window of the dynamics), we found that choosing appropriate learning rates ($\eta_{w}$ for inner optimization and $\eta_{J}$ for outer optimization) is essential for improving the model’s meta-generalization ability; neither rate should be too small. Additionally, we explored overparameterized meta-learners which have more hidden units than the meta-teachers, and showed the flexibility they offer in selecting learning rates. We also demonstrated that meta-learning algorithms can handle situations where the input-to-hidden weights of each task-specific teacher deviate from those of the meta-teacher. The proposed framework can also accommodate other learning scenarios, such as the effect of explicit regularization, other activation functions, and so on, where we briefly discussed in Appendix D and E. We believe that this work offers a solid framework and a wealth of examples for the statistical-mechanical analysis of meta-learning, with potential for extension to explore other interesting meta-learning scenarios.
Acknowledgements.
We thank David Saad for fruitful discussions. H.W. and B.L. acknowledged supports from the National Natural Science Foundation of China (Grant No. 12205066), the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515011775), the National Natural Science Foundation of China (Grant No. 12205066), the Shenzhen Start-Up Research Funds (Grant No. BL20230925) and the start-up funding from Harbin Institute of Technology, Shenzhen (Grant No. 20210134). C.T.Y. acknowledged supports from the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2214050004792), the National Natural Science Foundation of China (Grant No. 1217040938), the Shenzhen Municipal Science and Technology projects (Grant No. 202001093000117).
Appendix A Approximation of $\langle w_{i}w_{k}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}$
@@
When deriving the dynamical equations of the order parameter $R_{kn}$, we have made use of the de-correlation assumption of $\langle w_{i}w_{k}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}\approx\langle w_{i}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}\langle w_{k}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}$ in the limit of large $P$. Here, we briefly outline the justification for this approximation.
@@
Remind that different example inputs $\{\boldsymbol{\xi}^{\mu}\}$ in the same task are drawn independently from a certain distribution. Therefore, for $\mu^{\prime}\neq\mu$, $g(x_{i}^{\mu^{\prime}})g(y_{n}^{\mu^{\prime}})$ and $g(x_{k}^{\mu})g(y_{m}^{\mu})$ can be treated as independent random variables, and their averages can be performed separately
@@
Observe that the summation $\frac{1}{P^{2}}\sum_{\mu,\mu^{\prime}=1}^{P}(\cdots)$ in Eq. (44) have $P^{2}$ terms, while there are $P(P-1)$ terms satisfying $\mu^{\prime}\neq\mu$. In the limit where $P$ is large, then the summation $\frac{1}{P^{2}}\sum_{\mu,\mu^{\prime}=1}^{P}(\cdots)$ is dominated by the terms of $\mu^{\prime}\neq\mu$. It allows us to make the approximation
@@
Appendix B Self-Averaging of the Learning Dynamics in the Limit of Large $P$ and $V$
@@
When $V$ is large, the terms $\frac{1}{V}\sum_{\nu=1}^{V}g(y_{m}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})y_{n}^{t+\tau,\nu}$ and $\frac{1}{V}\sum_{\nu=1}^{V}g(x_{i}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})y_{n}^{t+\tau,\nu}$ appeared in Eq. (47) converges to their average values $I_{3}(k,n,m)$ and $I_{3}(k,n,i)$.
@@
We then further consider the limit where $\Delta t\to\infty$ but $\Delta t/N=\mathrm{d}\alpha\to 0$. In this limit, $\alpha$ plays the role of a continuous time variable, and the term $\frac{1}{N}\sum_{\tau=1}^{\Delta t}(\cdots)$ in the right hand side of Eq. (47) involves the summation of large number of random variables, where the randomness comes from $\{u_{m}^{t+\tau}\}$. We then invoke the law of large numbers and apply $\frac{1}{N}\sum_{\tau=1}^{\Delta t}(\cdots)\to\langle\cdots\rangle_{\boldsymbol{u}}$ in Eq. (47), which heuristically justifies the self-averaging property of dynamics of $R_{kn}$.
@@
Similar considerations apply to the dynamics of $Q_{kl}$, except that one needs to deal with averages of higher order terms like $\frac{1}{V}\sum_{\nu=1}^{V}g(y_{n}^{t+\tau,\nu})g(y_{m}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})g^{\prime}(x_{l}^{t+\tau,\nu})$, $\frac{1}{V}\sum_{\nu=1}^{V}g(y_{n}^{t+\tau,\nu})g(x_{i}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})g^{\prime}(x_{l}^{t+\tau,\nu})$ and $\frac{1}{V}\sum_{\nu=1}^{V}g(x_{i}^{t+\tau,\nu})g(x_{j}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})g^{\prime}(x_{l}^{t+\tau,\nu})$, which are usually subject to larger statistical fluctuations. Besides, these terms also have pre-factors that comprise of higher order terms of the learning rates like $\eta_{J}^{2}\eta_{w}^{2}$, $\eta_{J}^{2}\eta_{w}^{3}$ and $\eta_{J}^{2}\eta_{w}^{4}$. Therefore, higher values of learning rates can magnify the approximation or statistical errors when $V$ is finite, as we have seen in Fig. 2 of the main text. In Fig. 8 and Fig. 9, we show that for smaller values of $\eta_{w}$ and larger values of $V$, the theory provides a more accurate description of evolution of the order parameters $\{Q_{kl}\}$.
@@
Figure 8: Comparison of the theoretical and numerical results of $\{Q_{kl}\}$ for different values of $\eta_{w}$, where the experiment set-up is the same as the one in Sec. IV.1, except that a single run of the simulated dynamics is considered here. Solid lines representing experimental data and dashed lines corresponding to theoretical predictions. Panels (a) and (b) correspond to $\eta_{w}=4$. Panels (c) and (d) correspond to $\eta_{w}=1$.
@@
Figure 9: Comparison of the theoretical and numerical results of $\{Q_{kl}\}$ for different values of $P,V$, where the experiment set-up is the same as the one in Sec. IV.1, except that a single run of the simulated dynamics is considered here. Solid lines representing experimental data and dashed lines corresponding to theoretical predictions. Panels (a) and (d) correspond to $P=V=20$. Panels (b) and (e) correspond to $P=V=50$. Panels (c) and (f) correspond to $P=V=100$.
Appendix C Details on the Variability of Teacher Representation Mapping

### v2

@@ Appendix A More Details on Disorder Averages
where the detailed expressions can be found in Ref. [41].
Appendix B Approximation of $\langle w_{i}w_{k}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}$
@@
When deriving the dynamical equations of the order parameter $R_{kn}$, we have made use of the decorrelation assumption of $\langle w_{i}w_{k}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}\approx\langle w_{i}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}\langle w_{k}\rangle_{\{\boldsymbol{\xi}^{\mu}\}}$ in the limit of large $P$. Here we briefly outline the justification for this approximation.
@@
Recall that different example inputs $\{\boldsymbol{\xi}^{\mu}\}$ in the same task are drawn independently from a certain distribution. Therefore, for $\mu^{\prime}\neq\mu$, $g(x_{i}^{\mu^{\prime}})g(y_{n}^{\mu^{\prime}})$ and $g(x_{k}^{\mu})g(y_{m}^{\mu})$ can be treated as independent random variables, and their averages can be performed separately
@@
Observe that the summation $\frac{1}{P^{2}}\sum_{\mu,\mu^{\prime}=1}^{P}(\cdots)$ in Eq. (50) has $P^{2}$ terms, while there are $P(P-1)$ terms satisfying $\mu^{\prime}\neq\mu$. In the limit where $P$ is large, then the summation $\frac{1}{P^{2}}\sum_{\mu,\mu^{\prime}=1}^{P}(\cdots)$ is dominated by the terms of $\mu^{\prime}\neq\mu$. It allows us to make the approximation
@@
Appendix C Self-Averaging of the Learning Dynamics in the Limit of Large $P$ and $V$
@@
When $V$ is large, the terms $\frac{1}{V}\sum_{\nu=1}^{V}g(y_{m}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})y_{n}^{t+\tau,\nu}$ and $\frac{1}{V}\sum_{\nu=1}^{V}g(x_{i}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})y_{n}^{t+\tau,\nu}$ that appeared in Eq. (53) converge to their average values $I_{3}(k,n,m)$ and $I_{3}(k,n,i)$.
@@
We then further consider the limit where $\Delta t\to\infty$ but $\Delta t/N=\mathrm{d}\alpha\to 0$. In this limit, $\alpha$ plays the role of a continuous time variable, and the term $\frac{1}{N}\sum_{\tau=1}^{\Delta t}(\cdots)$ in the right-hand side of Eq. (53) involves the summation of large number of random variables, where the randomness comes from $\{u_{m}^{t+\tau}\}$. We then invoke the law of large numbers and apply $\frac{1}{N}\sum_{\tau=1}^{\Delta t}(\cdots)\to\langle\cdots\rangle_{\boldsymbol{u}}$ in Eq. (53), which heuristically justifies the self-averaging property of dynamics of $R_{kn}$.
@@
Similar considerations apply to the dynamics of $Q_{kl}$, except that one needs to deal with averages of higher order terms like $\frac{1}{V}\sum_{\nu=1}^{V}g(y_{n}^{t+\tau,\nu})g(y_{m}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})g^{\prime}(x_{l}^{t+\tau,\nu})$, $\frac{1}{V}\sum_{\nu=1}^{V}g(y_{n}^{t+\tau,\nu})g(x_{i}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})g^{\prime}(x_{l}^{t+\tau,\nu})$ and $\frac{1}{V}\sum_{\nu=1}^{V}g(x_{i}^{t+\tau,\nu})g(x_{j}^{t+\tau,\nu})g^{\prime}(x_{k}^{t+\tau,\nu})g^{\prime}(x_{l}^{t+\tau,\nu})$, which are usually subject to larger statistical fluctuations. As well, these terms also have prefactors that comprise of higher order terms of the learning rates like $\eta_{J}^{2}\eta_{w}^{2}$, $\eta_{J}^{2}\eta_{w}^{3}$ and $\eta_{J}^{2}\eta_{w}^{4}$. Therefore, higher values of learning rates can magnify the approximation or statistical errors when $V$ is finite, as we have seen in Fig. 2 . In Fig. 8 and Fig. 9, we show that for smaller values of $\eta_{w}$ and larger values of $V$, the theory provides a more accurate description of evolution of the order parameters $\{Q_{kl}\}$.
@@
Figure 8: Comparison of the theoretical and numerical results of $\{Q_{kl}\}$ for different values of $\eta_{w}$, where the experiment setup is the same as the one in Sec. IV.1, except that a single run of the simulated dynamics is considered here. Solid lines represent experimental data and dashed lines correspond to theoretical predictions. Panels (a) and (b) correspond to $\eta_{w}=4$. Panels (c) and (d) correspond to $\eta_{w}=1$.
@@
Figure 9: Comparison of the theoretical and numerical results of $\{Q_{kl}\}$ for different values of $P,V$, where the experiment setup is the same as the one in Sec. IV.1, except that a single run of the simulated dynamics is considered here. Solid lines represent experimental data and dashed lines correspond to theoretical predictions. Panels (a) and (d) correspond to $P=V=20$. Panels (b) and (e) correspond to $P=V=50$. Panels (c) and (f) correspond to $P=V=100$.
Appendix D Details on the Variability of Teacher Representation Mapping